The radius is 3 inches more than the height. The volume is 16π cubic inches. Find the dimensions of the right circular cylinder described.

Answer by Alan3354(69443)

The radius is 3 inches more than the height. The volume is 16pi cubic inches.
----
Vol = pi*r^2^h = 16pi
r^2h = 16
h = r-3
r^2*(r-3) = 16
r^3 - 3r^2 - 16 = 0
r = 4 is the only real root (by graphing)
h = 1